OpenStudy (anonymous):

Help on a few problems? Simplifying radicals. 1. 8√63x^5 2. √128x^5y^2 3. ∛32 4. ∛56x^14 I have the correct answers, but i have no idea how to get them? Radicals are not my strong suit..

OpenStudy (anonymous):

Anyone..

OpenStudy (anonymous):

\[ 8\sqrt{63x^5}=8(x^2)\sqrt{(9)(7)(x)}=8(3)(x^2)\sqrt{7x}=24x^2\sqrt{7x}\]\[\sqrt{128x^5y^2}=x^2y \sqrt{64(2)x}=8x^2y \sqrt{2x}\]\[\sqrt[3]{32}=\sqrt[3]{8(4)}=2\sqrt[3]{4}\]\[\sqrt[3]{56x^{14}}=x^4\sqrt[3]{8(7)x^2}=2x^4\sqrt[3]{7x^2}\]

OpenStudy (anonymous):

Is that right? Please check